# Estimation of AMOC transition probabilities using a machine learning based rare-event algorithm

### v1

@@ Abstract
The Atlantic Meridional Overturning Circulation (AMOC) is an important component of the global climate, known to be a tipping element, as it could collapse under global warming. The main objective of this study is to compute the probability that the AMOC collapses within a specified time window, using a rare-event algorithm called Trajectory-Adaptive Multilevel Splitting (TAMS). However, the efficiency and accuracy of TAMS depend on the choice of the score function. Although the definition of the optimal score function, called “committor function” is known, it is impossible in general to compute it a priori. Here, we combine TAMS with a Next-Generation Reservoir Computing technique that estimates the committor function from the data generated by the rare-event algorithm. We test this technique in a stochastic box model of the AMOC for which two types of transition exist, the so-called F(ast)-transitions and S(low)-transitions. Results for the F-transtions compare favorably with those in the literature where a physically-informed score function was used. We show that coupling a rare-event algorithm with machine learning allows for a correct estimation of transition probabilities, transition times, and even transition paths for a wide range of model parameters. We then extend these results to the more difficult problem of S-transitions in the same model. In both cases of F- and S-transitions, we also show how the Next-Generation Reservoir Computing technique can be interpreted to retrieve an analytical estimate of the committor function.
@@ 1 Introduction
Tipping elements (Lenton et al., 2008; Armstrong McKay et al., 2022) are subsystems of the climate system that may undergo rapid changes due to global warming. In particular, the Atlantic Meridional Overturning Circulation (AMOC) has been suggested to be in a bistable regime (Stommel, 1961). Since then, the AMOC collapse under freshwater forcing has been studied in various models (Rahmstorf, 1996), and was recently found at the higher end of the model complexity scale (van Westen and Dijkstra, 2023). Observations also suggest that the present-day AMOC is in such a bistable state (Bryden et al., 2011; Garzoli et al., 2013; Weijer et al., 2019; van Westen and Dijkstra, 2023). Considering the high impact of a collapse of the AMOC on the global climate, the probability that it occurs in the coming century should be assessed, even though this transition is a rare event.
@@
Due to its very low likelihood, this event cannot be studied by direct numerical simulations, even in conceptual models. From a theoretical point of view, Large Deviation Theory (Freidlin and Wentzell, 1998) is the natural framework for studying such events and has recently been applied (Soons et al., 2023) to a conceptual AMOC model. For this model (Wood et al., 2019), so-called optimal transition paths (instantons) could be determined, which show the most likely path for the AMOC collapse. In addition, the instantons provide the optimal freshwater forcing needed to effectively collapse the AMOC. However, due partly to the required mathematical assumptions, e.g. low noise, it is not guaranteed that such a method can be applied to fully-coupled climate models.
@@
On the other hand, Trajectory-Adaptive Multilevel Splitting (TAMS) (Lestang et al., 2018) was designed to estimate transition probabilities in fixed time between separate subsets of the phase space. More precisely, given two sets $A$ (e.g., present-day AMOC) and $B$ (e.g., collapsed AMOC), TAMS iteratively drives any trajectory starting in $A$ to reach $B$ within a chosen time, and returns the probability of occurrence of such a transition. This process relies on a score function that discards trajectories unlikely to transition, while cloning and resimulating promising ones without influencing the transition probability. This implies that the score function is key in ensuring the efficiency of TAMS: a bad choice of score function leads to increased variance in the probability estimates, which is also obtained at a (much) larger computational cost. This last point is especially important when applying TAMS to large-dimensional climate models, since their complexity makes simulated data scarce. This class of algorithms has already been successfully applied to climate problems, in particular extreme events, even in a climate model of full complexity (Ragone et al., 2018).
@@
In most cases, a prescribed score function was used in TAMS, although the optimal score function is known to be the committor function (Lestang et al., 2018; Cérou et al., 2019). It is a function of the phase space that associates to any point $\mathbf{x}$ the probability that trajectories starting from $\mathbf{x}$ reach $A$ before $B$. This function is so complex that it is impossible to compute exactly in general, even in simple dynamical systems (Lucente et al., 2022). Fortunately, there are several data-driven methods to estimate the committor function, and these have been compared by Jacques-Dumas et al. (2023): Analogues Markov Chain (Lucente et al., 2022), Feedforward Neural Networks, Next-Generation Reservoir Computing (Gauthier et al., 2021) and Dynamical Galerkin Approximation (Finkel et al., 2021). The Feedforward Neural networks were found to give the best score but the Next-Generation Reservoir Computing (RC) technique could achieve a similar score while requiring less computational time. The RC also presents several other advantages, such as the possibility to train it online and its interpretability.
@@
Here, we combine TAMS with RC to estimate transition probabilities while estimating the committor on-the-fly, using no more data than required by TAMS. We will apply this technique to the same conceptual AMOC model as in Castellana et al. (2019), for which there are two types of transitions: F(ast)-transitions during which the AMOC strength goes to zero due to a noisy freshwater forcing, and S(low)-transitions where the AMOC collapses to a different statistical equilibrium state. As TAMS was already applied to the F-transitions in Castellana et al. (2019) using a prescribed score function, these results serve as a comparison with the TAMS-RC method.
@@
In section 2, we present the conceptual AMOC model, the TAMS algorithm, the committor function, and the RC method with a focus on its combination with TAMS. Then, in section 3 we compare the results of the TAMS-RC method with those in Castellana et al. (2019). In Section 4, we extend the results of Castellana et al. (2019) to compute the probabilities of S-transitions. Finally, in section 5 we show how interpretable information can be extracted from trained neural networks, and in Section 6 we summarize and discuss the results and suggest possible future improvements.
@@ 2.1 The Cimatoribus-Castellana model
Conceptual models have been very useful for understanding possible multi-stable regimes of the AMOC (Dijkstra, 2023). Here, we use a conceptual AMOC model which was first introduced by Cimatoribus et al. (2014) and slightly modified by Castellana et al. (2019), with the configuration described in Jacques-Dumas et al. (2023). The Atlantic Ocean is represented by five boxes (Fig. 1). There is a northern Atlantic and a southern Atlantic box, respectively, labeled $n$ and $s$. The pycnocline layer is divided into two boxes: a tropical box (labeled $t$) and a southern tropical box (labeled $ts$, located south of $30^{\circ}$). Finally, there is a deep box (labeled $d$) that extends below the other four boxes, from the pycnocline depth to the ocean floor.
@@
In this model, the temperature of boxes n and ts is prescribed and fixed and the state vector consists simply of ($S_{t}$, $S_{t}s$, $S_{n}$, $S_{s}$, $S_{d}$, $D$), the salinity of the boxes and the pycnocline depth $D$. The volume transports between each box are represented by two wind-driven subtropical gyres $r_{s}$ and $r_{n}$, but most importantly by the three quantities: $q_{N}$, $q_{S}$ and $q_{U}$. $q_{N}$ corresponds to the downwelling that takes place in the Northern Atlantic, i.e. the AMOC strength. $q_{S}$ stands for the transport in the South Atlantic and is the difference between the wind-driven Ekman flow ($q_{Ek}$) and the eddy-induced volume transport ($q_{e}$). $q_{U}$ models the Ekman upwelling through the pycnocline.
@@
This model is forced by two separate freshwater fluxes. The first, $E_{s}$, is constant and symmetric from the box $t$ to the boxes $s$ and $n$. The second, $E_{a}$, is asymmetric and is applied from box $s$ to box $n$ and consists of a constant part and a white noise component, multiplied by an amplitude that can be controlled: $E_{a}(t)=\overline{E_{a}}(1+f_{\sigma}\zeta(t))$, where $\zeta(t)$ is a white noise with zero mean and unit variance. The choice of $(\overline{E_{a}},f_{\sigma})$ determines the behavior of the system. The complete equations of the model and the standard configuration of the parameters used are given in Jacques-Dumas et al. (2023) and are not repeated here.
@@
As explained in Castellana et al. (2019), for $\overline{E_{a}}\in[0.06,0.35]$, the AMOC in this model is in a bistable regime. Both its stable fixed points can be characterized in terms of volume transports. The first stable state has $q_{N}>q_{S}>q_{U}>0$; the downwelling in the North Atlantic is much stronger than the upwelling in the South Atlantic and the upwelling through the pycnocline, corresponding to a present-day-like AMOC circulation. The second stable state has $q_{N}=0\ \&\ q_{S}<0$, which means that the downwelling in the North Atlantic is shut down and the circulation in the Southern Hemisphere is reversed; this corresponds to a fully collapsed AMOC. For convenience, we will call these fixed points the “on” state and “off” state of the AMOC, respectively. A transition from the on-state to the off-state is called an S-transition and typically takes about $1,000$ years. Apart from S-transitions, F-transitions also occur in this model, during which the AMOC collapses to a state where $q_{N}=0\ \&\ q_{S}>0$. What happens in this case is that fast freshwater inputs due to noise may shut down the downwelling in the North Atlantic before the deep layer of the ocean can be disturbed, thus maintaining the upwelling in the South Atlantic. Since water accumulates above the pycnocline, this state is only transient, and either the AMOC recovers eventually or completely collapses. The typical timescale for such a temporary shutdown is about $100$ years.
@@ 2.2 Trajectory-Adaptive Multilevel Sampling (TAMS)
The general problem to determine the probability of either F- or S-transitions can be rephrased in more general terms as: given two areas of the phase space $A$ and $B$, what is the probability that a trajectory starting in $A$ reaches the zone $B$ before a certain time $T_{\mathrm{max}}$? For example, for an F-transition, $A$ would correspond to a neighborhood of the AMOC-on state, $B$ to $q_{N}=0$ and $T_{\mathrm{max}}=100$ years.
@@
This problem is hard, if not impossible, to treat with direct numerical simulation, especially in the case of rare events when there is a very low probability of observing a transition within a limited time. Splitting methods, on the other hand, are specially designed for this kind of tasks. In particular, the TAMS algorithm (Lestang et al., 2018) is designed to compute the probability of occurrence of rare events at a much lower cost than conventional Monte-Carlo simulations. At every iteration of TAMS, a score function $\Phi$ measures how close each simulated trajectory (among an ensemble) gets to $B$. The ones that are farthest (i.e., lowest score $\Phi$) are discarded and replaced by cloning and branching ”better” trajectories without affecting the transition probability. Thus, at each iteration, all trajectories reach a point that is increasingly close to $B$. When all trajectories have reached $B$, TAMS has converged and outputs an estimate of the transition probability as well as an ensemble of trajectories, each containing a transition.
The steps in the TAMS algorithm are described in Appendix A. Let $q$ be the real transition probability from $A$ to $B$. TAMS should be run $M>1$ times to obtain a distribution of its estimates $\hat{q}$. $\Phi$ is the most important quantity of TAMS, since it controls the deletion and mutation of trajectories and thus drives the trajectories towards $B$. Whatever score function $\Phi$ is used, TAMS is never biased (that is, $\mathbb{E}(\hat{q})=q$), but the variance of the distribution of $\hat{q}$ depends on $\Phi$ (Lestang et al., 2018).
The definition of the optimal score function of the TAMS algorithm, the committor function, is given in Appendix B. For example, for an F-transition in the AMOC model, the set $\mathcal{E}$ in Appendix B is the set $\{\mathbf{x}\in\mathbb{R}^{n}\ |\ q_{N}(\mathbf{x})\leq 0\}$ and the set $\mathcal{F}$ corresponds to $\{(T_{\mathrm{max}},\mathbf{x})\ |\ q_{N}(\mathbf{x})>0\}$. But since TAMS generates trajectories only until $T_{\mathrm{max}}$, the dependence of the committor on the mean passage time $\tau_{\mathcal{F}}$ is automatically implemented and the optimal score function for TAMS now reads (Lestang et al., 2018):
$\Phi(\mathbf{x})=\mathbb{P}(\underset{0\leq t\leq T_{\mathrm{max}}}{\mathrm{min}}q_{N}(\mathbf{X}(t))\leq 0\ |\ \mathbf{X}(0)=\mathbf{x}).$ (1)
Appendix A also gives an expression for the optimal variance $\sigma^{2}_{\mathrm{id}}$ of $\hat{q}$ (when the score function is the committor function).
@@
2.3 Estimating the score function using machine learning for the AMOC model
Our objective is to approach $\Phi$ using the least amount of data possible, that is, by learning this function while running TAMS, using only the data generated by TAMS. In Jacques-Dumas et al. (2023) we studied four data-driven methods that can estimate the committor. We compared their performance in terms of accuracy and computation time and found that the RC method (Gauthier et al., 2021) stood out for complex problems. The RC method (shortly described in Appendix C) is used here with the exact same setup as in Jacques-Dumas et al. (2023).
@@ 2.3 Estimating the score function using machine learning for the AMOC model
The combination of TAMS and RC does not pose any major problem: in the code for TAMS, the score function is replaced by the application of the RC to any trajectory where the committor must be computed. What requires more attention are the initialization and training process of the RC. The RC is initialized with $1,000$ trajectories generated for $(\overline{E_{a}}=0.35,f_{\sigma}=0.5)$, where the transition probability is $1$, which makes it easy to obtain transitions, allowing the RC to grasp the structure of the phase space. Due to the lack of data in the low transition probability regime, it ensures that the RC can be initialized at all. During the TAMS run, all discarded trajectories and all trajectories that reach the target set are used to retrain the RC at each iteration. When TAMS finishes, this trained RC is used as the initial score function of TAMS for the parameters $(\overline{E_{a}}=0.35,f_{\sigma}=0.5-\Delta_{f})$ and trained again during the run of TAMS. Setting $\Delta_{f}=0.005$, this process is repeated until $f_{\sigma}=0.005$. For every value of $f_{\sigma}$, the RC trained for $\overline{E_{a}}=0.35$Sv is used as the initial score function for TAMS with $\overline{E_{a}}=0.35\mathrm{Sv}-\Delta_{E_{a}}$ (with $\Delta_{E_{a}}=0.003$) and retrained as explained above. In this way, we cover the entire bistability regime, iteratively adapting the RC to smaller values of $\overline{E_{a}}$. This process is a form of transfer learning: information learned in an easier case (larger noise of forcing leading to a higher transition probability) is transferred to a more difficult case (i.e., lower transition probability) to adapt the neural network faster to this new situation.
Therefore, the RC is constantly retrained during each TAMS run, which means that the quality of the learned score function is highly dependent on the trajectories obtained from TAMS. Since they are subject to noise and randomly cloned, the learned score changes at each TAMS run. To compensate for this effect, we run the combination TAMS-RC $M$ times for each couple of parameters, thus training $M$ independent RC. Each RC is trained using only the trajectories obtained in its own run. This situation is more realistic in the case where we are working with a very large model and we cannot afford to run TAMS many times to optimize the training of the RC.

### v3

@@ Abstract
The Atlantic Meridional Overturning Circulation (AMOC) is an important component of the global climate, known to be a tipping element, as it could collapse under global warming. The main objective of this study is to compute the probability that the AMOC collapses within a specified time window, using a rare-event algorithm called Trajectory-Adaptive Multilevel Splitting (TAMS). However, the efficiency and accuracy of TAMS depend on the choice of the score function. Although the definition of the optimal score function, called “committor function” is known, it is impossible in general to compute it a priori. Here, we combine TAMS with a Next-Generation Reservoir Computing technique that estimates the committor function from the data generated by the rare-event algorithm. We test this technique in a stochastic box model of the AMOC for which two types of transition exist, the so-called F(ast)-transitions and S(low)-transitions. Results for the F-transtions compare favorably with those in the literature where a physically-informed score function was used. We show that coupling a rare-event algorithm with machine learning allows for a correct estimation of transition probabilities, transition times, and even transition paths for a wide range of model parameters. We then extend these results to the more difficult problem of S-transitions in the same model. In both cases of F-transitions and S-transitions, we also show how the Next-Generation Reservoir Computing technique can be interpreted to retrieve an analytical estimate of the committor function.
@@ 1 Introduction
Tipping elements (Lenton et al., 2008; Armstrong McKay et al., 2022) are subsystems of the climate system that may undergo rapid changes due to global warming. In particular, the Atlantic Meridional Overturning Circulation (AMOC) has been suggested to be in a bistable regime (Stommel, 1961). The AMOC collapse under freshwater forcing has since been studied in various models (Rahmstorf, 1996) and was recently found at the higher end of the model complexity scale (van Westen and Dijkstra, 2023). Observations also suggest that the present-day AMOC is in a bistable state (Bryden et al., 2011; Garzoli et al., 2013; Weijer et al., 2019; van Westen and Dijkstra, 2023). An AMOC collapse would have a global impact, for example a decrease of the surface temperature over the Northern Hemisphere and especially Europe (Jackson et al., 2015; van Westen et al., 2024; Cini et al., 2024) and a change of rainfall patterns over tropical regions (Jackson et al., 2015; Parsons et al., 2014; Chang et al., 2008). Although doubts exist about the possibility of this collapse and the scale of its consequences (Roquet and Wunsch, 2022), the potential impact of such an event calls for an assessment of its probability of occurrence before $2100$.
@@
Due to its very low likelihood, this event cannot be studied by direct numerical simulations. From a theoretical point of view, Large Deviation Theory (Freidlin and Wentzell, 1998) is the natural framework to study it and has recently been applied to a conceptual AMOC model (Soons et al., 2023) for which the instanton, or the most likely path to the collapse, has been computed. This method also provides the optimal freshwater forcing needed to effectively collapse the AMOC. However, due partly to the required mathematical assumptions, e.g. low noise, this method is not guaranteed to be applicable to fully-coupled climate models.
@@
On the other hand, Trajectory-Adaptive Multilevel Splitting (TAMS) (Lestang et al., 2018) was designed to estimate the probability of any transition in fixed time. Given two sets $A$ (e.g., present-day AMOC) and $B$ (e.g., collapsed AMOC), TAMS drives any trajectory starting in $A$ towards $B$ within a chosen time. This is done with a score function that discards trajectories unlikely to transition, while cloning and resimulating promising ones. This function is key in the efficiency of TAMS: a bad score function leads to increased variance in the probability estimates and makes them more costly to compute. This is especially important in large-dimensional climate models, whose complexity makes simulated data scarce. Similar algorithms have already been successfully applied to climate problems, in particular extreme events, even in a climate model of full complexity (Ragone et al., 2018).
@@
The optimal score function of TAMS is the committor function (Lestang et al., 2018; Cérou et al., 2019). It associates to any state $\mathbf{x}$ in the phase space the probability that trajectories starting from $\mathbf{x}$ reach $A$ before $B$. However, computing it exactly is already impossible in simple dynamical systems (Lucente et al., 2022). Fortunately, several data-driven methods exist to estimate this function: Analogues Markov Chain (Lucente et al., 2022), Feedforward Neural Networks, Next-Generation Reservoir Computing (Gauthier et al., 2021) and Dynamical Galerkin Approximation (Finkel et al., 2021). The Feedforward Neural networks were found to give the best committor estimate (Jacques-Dumas et al., 2023) but the Next-Generation Reservoir Computing (RC) technique achieved a similar performance while requiring less computational time. Other advantages of the RC include the possibility to train it online and its interpretability.
@@
The purpose of Jacques-Dumas et al. (2023) was to develop a benchmark for several committor estimation techniques to determine which is best fit for combination with TAMS. Here, we combine TAMS with one of them (RC) to compute transition probabilities while estimating the committor on-the-fly, using no more data than required by TAMS. We apply this technique to the same conceptual AMOC model as in Castellana et al. (2019), for which there are two types of transitions: F(ast)-transitions, noise-driven transitions during which random freshwater forcings drive the AMOC strength to zero, and S(low)-transitions where the AMOC shifts to another statistical equilibrium. This model is interesting for our study because these two kinds of collapse develop on different timescales. Moreover, Castellana et al. (2019) already used TAMS to compute the probability of occurrence of F-transitions. These results were obtained using a prescribed score function and are compared to the TAMS-RC method.
@@
In section 2, we present the conceptual AMOC model, the TAMS algorithm, the committor function, and the RC method with a focus on its combination with TAMS. Then, in section 3 we compare the results of the TAMS-RC method with those in Castellana et al. (2019). In Section 4, we extend the results of Castellana et al. (2019) to compute the probabilities of S-transitions. Finally, in section 5 we show how interpretable information can be extracted from the RC, and in Section 6 we summarize and discuss the results and suggest possible future improvements.
@@ 2.1 The Cimatoribus-Castellana model
Conceptual models have been very useful for understanding possible multi-stable regimes of the AMOC (Dijkstra, 2023). Here, we use a conceptual AMOC model introduced by Cimatoribus et al. (2014) with the configuration described in Castellana et al. (2019). The Atlantic Ocean is represented by five boxes (Fig. 1). The northern Atlantic and southern Atlantic boxes are respectively labeled $n$ and $s$. The pycnocline layer is divided into two boxes: a tropical box (labeled $t$) and a southern tropical box (labeled $ts$, located south of $30^{\circ}$). Finally, there is a deep box (labeled $d$) extending below the other boxes, from the pycnocline to the ocean floor.
@@
The temperature of all boxes is prescribed and fixed so that the state vector $\mathbf{x}=(S_{t},\ S_{ts},\ S_{n},\ S_{s},\ S_{d},\ D)$ consists of the salinity of the boxes and the pycnocline depth $D$. The volume transports between each box are represented by two wind-driven subtropical gyres $r_{s}$ and $r_{n}$ and three quantities: $q_{N}$, $q_{S}$ and $q_{U}$. The quantity $q_{N}$ corresponds to the downwelling taking place in the Northern Atlantic, i.e. the AMOC strength. The variable $q_{S}$ stands for the transport in the Southern Atlantic and is the difference between the wind-driven Ekman flow ($q_{Ek}$) and the eddy-induced volume transport ($q_{e}$). The quantity $q_{U}$ represents the Ekman upwelling through the pycnocline.
@@
This model is forced by two freshwater fluxes. The first, $E_{s}$, is constant and symmetric from the box $t$ to the boxes $s$ and $n$. The second, $E_{a}$, is asymmetric and applied from box $s$ to box $n$. It models a physical asymmetry between both boxes that controls the strength of the AMOC while keeping the salinity budget closed. The asymmetric freshwater flux reads: $E_{a}(t)=\overline{E_{a}}(1+f_{\sigma}\zeta(t))$ where $\overline{E_{a}}$ and $f_{\sigma}$ are constant and $\zeta(t)$ is a white noise with zero mean and unit variance. The choice of $(\overline{E_{a}},f_{\sigma})$ thus determines the (stochastic) behavior of the system. The complete equations of the model and the standard configuration of the parameters are given in Jacques-Dumas et al. (2023) and not repeated here.
@@
In this model, the AMOC is in a bistable regime for $\overline{E_{a}}\in[0.06,0.35]\mathrm{\ Sv}$ ($1\mathrm{\ Sv}=10^{6}\mathrm{\ m}^{3}$). Both its stable fixed states can be characterized in terms of volume transports. The first stable state has $q_{N}>q_{S}>q_{U}>0$; the downwelling in the Northern Atlantic is much stronger than the upwelling in the Southern Atlantic and through the pycnocline, corresponding to a present-day-like AMOC circulation. The second stable state has $q_{N}=0$ and $q_{S}<0$, such that the downwelling in the Northern Atlantic is shut down and the circulation in the Southern Hemisphere is reversed; this corresponds to a fully collapsed AMOC (see dotted red arrows on Fig. 1). We call these fixed states the “on” state and “off” state of the AMOC, respectively. A transition from the on-state to the off-state is called an S-transition and typically takes about $1,000$ years. Apart from S-transitions, F-transitions also occur in this model, during which the AMOC collapses to a state where $q_{N}=0$ and $q_{S}>0$. In that case, fast noisy freshwater inputs shut down the downwelling in the Northern Atlantic before the deep layer of the ocean can be disturbed, thus maintaining the upwelling in the Southern Atlantic. Since water accumulates above the pycnocline, this state is only transient and the AMOC either recovers or completely collapses (leading to an S-transition). The typical timescale for such a temporary shutdown is about $100$ years.
@@ 2.2 Trajectory-Adaptive Multilevel Sampling (TAMS)
The problem of determining the probability of either F-transitions or S-transitions can be rephrased as: given two areas of the phase space $A$ and $B$, what is the probability that a trajectory starting in $A$ reaches $B$ before a time $T_{\mathrm{max}}$? For an F-transition, $A$ corresponds to a neighborhood of the AMOC-on state, $B$ to $q_{N}=0$ and $T_{\mathrm{max}}=100$ years.
@@
This problem is impossible to treat with direct numerical simulation, especially in the case of rare events due to the low transition probability within a limited time. However, splitting methods are designed for this kind of task. In particular, the TAMS algorithm (Lestang et al., 2018) computes the probability of occurrence of rare events at a much lower cost than Monte-Carlo simulations. This algorithm is detailed in Appendix A. At every iteration of TAMS, a score function $\Phi$ measures how close each simulated trajectory (among an ensemble) gets to $B$. The farthest ones (i.e., lowest score $\Phi$) are discarded and replaced by cloning and branching ”better” trajectories. Thus, at each iteration, all trajectories get closer to $B$. TAMS is said to converge when almost all trajectories have reached $B$ (see Appendix A) and it outputs an estimate of the transition probability as well as an ensemble of trajectories, each containing a transition. TAMS thus efficiently generates an ensemble of reactive trajectories while allowing to trace back to the original transition probability of the model.
Let $\alpha$ be the transition probability from $A$ to $B$ within $T_{\mathrm{max}}$. Running TAMS $M>1$ times yields a distribution of its estimates $\hat{\alpha}$. Whatever score function $\Phi$ is used, $\mathbb{E}(\hat{\alpha})=\alpha$ in the limit of infinite $M$, but the variance of the distribution of $\hat{\alpha}$ depends on $\Phi$ (Lestang et al., 2018). The optimal score function for TAMS, which minimizes this variance, is the committor function (see Lestang et al. (2018) and Lucente et al. (2022) for its definition and properties). Appendix A also gives an expression for the optimal variance $\sigma^{2}_{\mathrm{id}}$ of $\hat{\alpha}$ when the committor is used as score function.
@@
2.3 Estimating the committor function using machine learning for the AMOC model
We aim at approaching the committor function while using only the data generated by TAMS. In Jacques-Dumas et al. (2023) we studied four data-driven methods that can perform this task. In terms of accuracy and computation time, we found Next-Generation Reservoir Computing (RC) (Gauthier et al., 2021) to stand out. This method maps the states of a dynamical system onto a larger-dimensional phase space using a family of nonlinear functions. Here, we take monomials of degree $p=4$ that are products of powers of $S_{n},S_{ts},D$ and the time $t$. Our implementation of this method is presented in Appendix B and with more details in Jacques-Dumas et al. (2023).
@@ 2.3 Estimating the committor function using machine learning for the AMOC model
Combining TAMS and RC poses the problems of initialization and the training process, since enough data is needed to obtain reliable predictions. The training of the RC is initialized with $1,000$ trajectories generated for $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma}=0.5)$, where the transition probability is $1$, which makes it easy to obtain transitions. For each TAMS run with parameters $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma})$, all discarded trajectories and all trajectories reaching the target set are used to update the RC at each iteration (see Appendix B). When TAMS converges, this trained RC is used as the initial score function of TAMS for parameters $(\overline{E_{a}}=0.35\mathrm{\ Sv},f_{\sigma}-\Delta_{f})$ with $\Delta_{f}=0.005$ and this process is repeated until $f_{\sigma}=0.005$. For every value of $f_{\sigma}$, the RC trained for $\overline{E_{a}}=0.35\mathrm{\ Sv}$ is used as the initial score function for TAMS with $\overline{E_{a}}=0.35\mathrm{\ Sv}-\Delta_{E_{a}}$ (with $\Delta_{E_{a}}=0.003\mathrm{\ Sv}$) and retrained as explained above. In this way, the RC adapts iteratively to parameters for which transitions are increasingly unlikely. This process is a form of transfer learning: information learned in an easier case (i.e., larger noise or forcing leading to a large transition probability) is transferred to a more difficult case (i.e., lower transition probability) to adapt the RC faster to this new situation. Each RC is trained using only the trajectories obtained in its own run, making them all independent. The transfer learning could also be performed by first looping over $\overline{E_{A}}$, then $f_{\sigma}$. We tested this alternative setup and found that the results hardly changed overall.